OpenStudy (anonymous):

is (2x)^2 the same as 2x^2

OpenStudy (raden):

no, it means (2x)^2 = 2^2 * x^2 = 4x^2

OpenStudy (raden):

remmber one of property of exponent : (ab)^m = a^m * b^m

OpenStudy (anonymous):

Thanks!